# Optimal Control of an Epidemic with Intervention Design

### v2

@@ 3.5.1 Adjoint system and Shadow Prices
$\dot{\lambda}_{i}$ $=-\left[c_{H}h+c_{NH}+2\kappa(i-I_{\max})_{+}\right]e^{-\delta t}+(\lambda_{s}-\lambda_{e})(\beta-h)s+\gamma\lambda_{i}.$ (10)
Economically, $\lambda_{s}(t)$ represents the marginal shadow cost of susceptibility (the reduction in future objective value achievable by removing an individual from the susceptible compartment), while $\lambda_{i}(t)$ represents the marginal social cost of infection. The term $2\kappa(i-I_{\max})_{+}$ acts as a soft barrier, signaling high costs when capacity is breached.
@@ 3.5.2 Switching functions and Optimality
3.5.3 Singular Regimes
Singular arcs occur if a switching function vanishes on a nontrivial interval. We analyze singular regimes as candidate solutions consistent with binding capacity constraints. A critical regime is the boundary-maintenance arc, where the planner keeps the system at capacity:
@@ 3.5.3 Singular Regimes
$h_{\mathrm{bm}}(t)=\beta-\frac{\gamma}{s(t)}\quad\iff\quad\mathcal{R}_{\mathrm{eff}}(t):=\frac{\beta-h_{\mathrm{bm}}(t)}{\gamma}s(t)=1.$
@@ 4.2 Optimal Control Dynamics and Shadow Prices
The economic rationale for this policy is revealed by the adjoint variables, or shadow prices, in Figure 4. The shadow price of infection, $\lambda_{i}(t)$, represents the marginal cost of an additional infected individual; it begins at a peak value of approximately 20.0 and decays as the epidemic is brought under control. The trajectory of $\lambda_{s}(t)$ represents the marginal value of preserving susceptible mass to prevent downstream costs.
@@ 5 Conclusion and Future Directions
Analytically, we characterized the optimal intervention strategy using the Pontryagin Maximum Principle, identifying singular arcs where the optimal policy maintains the epidemic trajectory precisely at the healthcare capacity limit. This "boundary-maintenance" regime represents a critical operational insight often missed by purely descriptive models. Furthermore, our derivation of a time-free representation of the dynamics and the use of the Lambert $W$ function provided sharp bounds on the final epidemic size under maximal suppression.

### v3

@@ 3.5.1 Adjoint system and Shadow Prices
$\dot{\lambda}_{i}$ $=-\left[c_{H}h+c_{NH}+2\kappa\max\{0,\,i-I_{\max}\}\right]e^{-\delta t}+(\lambda_{s}-\lambda_{e})(\beta-h)s+\gamma\lambda_{i}.$ (10)
Economically, $\lambda_{s}(t)$ represents the marginal shadow cost of susceptibility (the reduction in future objective value achievable by removing an individual from the susceptible compartment), while $\lambda_{i}(t)$ represents the marginal social cost of infection. The term $2\kappa\max\{0,\,i-I_{\max}\}$ acts as a soft barrier, signaling high costs when capacity is breached.
@@ 3.5.2 Switching functions and Optimality
3.5.3 Boundary-Maintenance Arcs
In optimal control theory, a singular arc refers to an interval on which a switching function vanishes identically, so that the optimal control is not determined by the bang-bang law alone. A distinct but related phenomenon is the boundary-maintenance arc, which arises when the optimal trajectory reaches a state (phase) constraint and is held at its boundary. We analyze this latter regime, in which the infection level is maintained at the capacity limit.
@@ 3.5.3 Boundary-Maintenance Arcs
$h_{\mathrm{bm}}(t)=\beta-\frac{\gamma}{s(t)}\quad\iff\quad\mathcal{R}_{\mathrm{eff}}(t):=\frac{\bigl(\beta-h_{\mathrm{bm}}(t)\bigr)}{\gamma}s(t)=1.$
Here $\mathcal{R}_{\mathrm{eff}}(t)$ is the effective reproduction number at time $t$: the expected number of secondary infections produced by a single infected individual given the current susceptible fraction $s(t)$ and suppression level $h_{\mathrm{bm}}(t)$. The condition $\mathcal{R}_{\mathrm{eff}}(t)=1$ characterizes the boundary-maintenance arc as the regime in which the epidemic is held in exact equilibrium, neither growing nor declining.
@@ 4.2 Optimal Control Dynamics and Shadow Prices
The economic rationale for this policy is revealed by the adjoint variables, or shadow prices, in Figure 4. The shadow price of infection, $\lambda_{i}(t)$, represents the marginal cost of an additional infected individual; it begins at a peak value of approximately 20.0 and decays as the epidemic is brought under control. The trajectory of $\lambda_{s}(t)$ represents the marginal value of preserving the susceptible mass to avoid downstream costs. It should be noted that the optimal control simulation presented in Figure 3 uses $t_{\mathrm{delay}}=0$, so that vaccination is available from the outset. The impact of positive delays on total cost is captured in the sensitivity analysis of Section 4.3 (Figure 6), which demonstrates that increasing $t_{\mathrm{delay}}$ substantially raises the median total cost. A detailed comparison of optimal trajectories with and without the delay is an important direction for future numerical investigation.
@@ 5 Conclusion and Future Directions
Analytically, we characterized the optimal intervention strategy using the Pontryagin Maximum Principle, identifying boundary-maintenance arcs where the optimal policy maintains the epidemic trajectory precisely at the healthcare capacity limit. This "boundary-maintenance" regime represents a critical operational insight often missed by purely descriptive models. Furthermore, our derivation of a time-free representation of the dynamics and the use of the Lambert $W$ function provided sharp bounds on the final epidemic size under maximal suppression.